OpenStudy (ilovehorses14):

The capacity of one container is 10 kiloliters. The capacity of another container is 20% greater than the first container. What is the actual capacity of the second container? A. 10 kl B. 12 kl C. 15 kl D. 20 kl

OpenStudy (michele_laino):

the requested capacity, is given by the subsequent computation: \[\huge 10 + \left( {\frac{{20}}{{100}} \times 10} \right) = ...?\]